OpenStudy (anonymous):

An arrow is shot into the air from a building that is 32 ft. high. The height of the arrow after t seconds is given by the formula h(t)= -16t²+48t+32=0 At what times will the arrow be at a height of 64 ft?

OpenStudy (anonymous):

Set the equation equal to 64 and solve.

OpenStudy (anonymous):

first, use the derivative to find out the function for velocity

OpenStudy (anonymous):

This is a superfluous step. The function gives height. You are solving for height. No need to differentiate. Just a quadratic function.

OpenStudy (anonymous):

never mind, I thought too much

OpenStudy (anonymous):

You are solving for time in a function of height...is what that should read. Apologies.

OpenStudy (anonymous):

so i should still set the equation = to 64 and then what will i do?

OpenStudy (anonymous):

Solve the equation for t.

OpenStudy (anonymous):

so i should subtract the 64 or should i factor the formula and then set it = to 64 and solve?

OpenStudy (anonymous):

You would have: \[-16t^2 + 48t + 32 = 64\] You can certainly factor a -16 out of the left

OpenStudy (anonymous):

Then what do you have?

OpenStudy (anonymous):

-16(t²-48t-32)=64?

OpenStudy (anonymous):

No, 16 will also go into 48 and 32. This is not factored all of the way.

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

-16(t²-3t-2)=64?

OpenStudy (anonymous):

Yep, now divide out the -16.

OpenStudy (anonymous):

=-4 and then i just factor it and then divide?

OpenStudy (anonymous):

t²-3t-2=-4 what do i do now?

OpenStudy (anonymous):

Yes, well you would factor this by completing the square. Then what do you have?

OpenStudy (anonymous):

Hint: You can get REALLY close with (t - 1)(t - 2)

OpenStudy (anonymous):

can i add the move the -2 from the left to the right to get =-4 +2 then i move that -2 to the left to get t²-3t+2?

OpenStudy (anonymous):

Well, what happens when you distribute out (t-1)(t-2) ?

OpenStudy (anonymous):

it equals t²-3t+2

OpenStudy (anonymous):

Right, but that doesn't equate to our original left hand side. It's greater by 4. So how do we make the +2 into a -2 ?

OpenStudy (anonymous):

to get the answer i have to complete the square right?

OpenStudy (anonymous):

Yes and that's what we're doing. To change the last term (+2) into (-2) we have to subtract 4. So you have: \[(t-1)(t-2) - 4 = -4\] Now add the 4 to both sides, what do you have?

OpenStudy (anonymous):

(t-1)(t-2)=0? i'm confused on how it went from t²-3t-2=-4 to get that

OpenStudy (anonymous):

To complete the square we realize that we can get the first two terms easily, but the last term doesn't come out right: (t-1)(t-2) = x^2 - 3x + 2 However, our original expression was x^2 - 3x - 2 Close but no cigar. We need to turn that +2 to a -2. To do this we subtract 4. Therefore we have the nice term: (t-1)(t-2)-4 = -4

OpenStudy (anonymous):

We're just making the left hand side look different. Making it easier to work with.

OpenStudy (anonymous):

where did you get the subtraction of -4 though is where i'm confused on

OpenStudy (anonymous):

Because when we foil out the expression (t-1)(t-2) and subtract 4, we get our ORIGINAL expression back.

OpenStudy (anonymous):

Ok, FOIL out (t-1)(t-2)

OpenStudy (anonymous):

so (t-1)(t-2) is = to t²-3t+2 and what you want me to do is factor out of t²-3t-2, (t-1)(t-2)?

OpenStudy (anonymous):

Well, follow me here. If (t-1)(t-2) = t^2 - 3t + 2, what happens when we subtract 4?

OpenStudy (anonymous):

if we subtract 4 it turns into t^2-3t-2

OpenStudy (anonymous):

Thats our ORIGINAL expression isnt it? :]

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

Right, if we didnt subtract out the 4, then we have just borked the entire equation and the solutions would be incorrect. The key in this is making it LOOK different while still meaning the same thing. We are ignoring the right hand side, btw...

OpenStudy (anonymous):

If we didn't complete the square we would have to use the quadratic formula which is cumbersome.

OpenStudy (anonymous):

\[x^2 - 3x - 2 = (x-1)(x-2)-4\]

OpenStudy (anonymous):

oh okay i see now

OpenStudy (anonymous):

Right, so now add the 4 and you simply have (t-1)(t-2) = 0

OpenStudy (anonymous):

Now what are the solutions?

OpenStudy (anonymous):

t=1 or t=2

OpenStudy (anonymous):

t = 1 AND t = 2. At 1 second, it's at 64ft on the way UP, at 2 seconds its at 64ft on the way DOWN.

OpenStudy (anonymous):

OH i see now and i fully get it. thank you so much

OpenStudy (anonymous):

No problem.